please help me with this!
x squared -9x+14

Answer by checkley71(8403)

X~2-9X+14=0 OR (X-7)(X-2)=0 OR X-7=0 OR X=7 & X-2=0 OR X=2
PROOF 7~2-9*7+14=0 OR 49-63+14=0 OR 63-63=0 OR 0=0
& 2~2-9*2+14=0 OR 4-18+14=0 OR 18-18=0 OR 0=0